# Motion-Aware Transformer for Multi-Object Tracking

### v1

@@ 1 Introduction
Multi-object tracking (MOT) aims to detect objects and maintain their identities across frames in a video. Recent advances in query-based detection have reshaped this task. The introduction of DETR Carion et al. (2020) marked a turning point, as it replaced handcrafted post-processing with an end-to-end formulation. Through attention mechanisms and Hungarian matching, DETR aligns queries with ground-truth objects, while unmatched queries are discarded as “no object.” This query-based design allows the network to propagate information across frames, making end-to-end training feasible and effective.
@@
Building on this paradigm, TrackFormer Meinhardt et al. (2022) extended DETR to MOT by propagating detection queries forward as track queries, and concatenating them with newly initialized detection queries for subsequent frames. This formulation enabled consistent learning of track queries without manual association rules. Later, MOTR Zeng et al. (2022) introduced a Temporal Aggregation Network (TAN) to update track queries between frames, alleviating issues of overfitting and improving performance on challenging datasets such as DanceTrack. These advances have established query propagation as a dominant strategy in end-to-end MOT.
@@
However, a key limitation remains: existing approaches typically process detection and track queries simultaneously within a single Transformer Decoder layer. This design introduces what we call query collisions. Track queries are required to follow the same object consistently across frames, whereas detection queries are reassigned at each frame through Hungarian matching. When a track query drifts from its ground-truth location, Hungarian matching may assign it to a different object that happens to be closer, resulting in identity switches and unstable training. At the same time, detection queries suffer from noisy gradients caused by drifting track queries, further reducing association performance. Figure 1(b) illustrates a typical collision scenario: a track query with low overlap with its true object is instead matched to a nearby distractor, while the correct detection query is left unmatched. Such conflicts degrade both training and inference performance, particularly in crowded or fast-moving scenes.
Figure 1: (a) Distribution of track query distances ($1-\text{IoU}$) for different methods. MATR (purple) concentrates track queries closer to ground truth compared to MOTR (cyan). (b) Illustration of query collisions in MOTR: a drifting track query (red) is wrongly assigned to object B under Hungarian matching, while MATR predicts the future position and keeps the track consistent.
@@
To overcome these limitations, we propose the Motion-Aware Transformer (MATR). MATR explicitly predicts the motion of track queries across frames, updating both their features and positional embeddings before they enter the Transformer Decoder. By anticipating object movements, MATR reduces the gap between track queries and their ground-truth targets, thereby minimizing query collisions and aligning the training process more closely with inference behavior. This motion-aware update provides consistent optimization for both detection and tracking queries, leading to substantial gains in association accuracy without adding complex components. As shown in Figure 1(a), MATR significantly lowers the distance between queries and their assigned objects compared to MOTR, confirming the effectiveness of motion-aware prediction.
@@
In addition to introducing MATR, we strengthen the MOTR baseline with improved training strategies and updated DETR-family components such as DAB-DETR Liu et al. (2022). This provides a stronger foundation for comparison and ensures that the observed improvements stem from the proposed motion-aware design. Comprehensive experiments across three benchmarks validate our approach. On the DanceTrack dataset, MATR improves HOTA by more than 9 points over MOTR without additional data, and reaches a state-of-the-art HOTA of 71.3 when supplementary data are included. On SportsMOT, MATR achieves a HOTA of 72.2, establishing a new state of the art without relying on extra datasets. Finally, in multi-category tracking on BDD100k, MATR obtains 54.7 mTETA and 41.6 mHOTA, surpassing previous methods under the same training conditions.
@@
In summary, MATR addresses the overlooked issue of query collisions in DETR-style MOT by introducing explicit motion prediction. This simple yet effective design improves association accuracy across diverse scenarios, leading to state-of-the-art performance while maintaining the elegance and efficiency of an end-to-end Transformer-based tracker.
@@ 2 Related Work
2.1 Tracking-by-Detection
Tracking-by-Detection remains the most common paradigm for multi-object tracking. These methods typically adopt a two-step pipeline: first detecting bounding boxes in each frame, then associating detections across frames to construct object trajectories. Performance is therefore heavily dependent on the accuracy of the detection stage.
SORT Bewley et al. (2016) exemplifies this approach by using the Hungarian algorithm (Kuhn, 1955) for association and a Kalman filter (Welch & Bishop, 1995) to predict object positions. Detections are matched to predicted boxes using intersection-over-union (IoU) scores. DeepSORT Wojke et al. (2017) extends SORT by incorporating appearance features extracted by a learned embedding network, which are compared via cosine distance to improve association robustness. Several methods integrate detection and re-identification more tightly. JDE Wang et al. (2020), FairMOT Zhang et al. (2021), and Unicorn Yan et al. (2022) jointly train detection and appearance embeddings in a unified framework.
ByteTrack Zhang et al. (2022), based on a YOLOX detector, achieves state-of-the-art performance by considering both high- and low-confidence detections, improving recall during association. BoT-SORT Aharon et al. (2022) further enhances SORT by refining the Kalman filter state, introducing camera-motion compensation, and integrating Re-ID features. Recent efforts incorporate transformers into the association stage. TransMOT Chu et al. (2023) and GTR Zhou et al. (2022) use spatial-temporal transformers to model interactions between instances and aggregate historical context. OC-SORT Cao et al. (2023) improves upon SORT by re-identifying and rehabilitating lost tracks, further demonstrating the importance of temporal reasoning in tracking-by-detection systems.
2.2 Query Propagation
The emergence of DETR-family models has popularized query propagation as an alternative, end-to-end approach for multi-object tracking. These methods maintain query embeddings across frames to track individual objects, eliminating the need for handcrafted post-processing.
@@ 2.2 Query Propagation
Tracktor++ Bergmann et al. (2019) pioneered this idea by leveraging an R-CNN Girshick et al. (2014) regression head to re-localize bounding boxes across frames. TrackFormer Meinhardt et al. (2022) extended Deformable DETR Zhu et al. (2021) to propagate detection queries as track queries, while MOTR Zeng et al. (2022) introduced a Temporal Aggregation Network (TAN) to continuously update track queries, achieving strong results on DanceTrack. MeMOT Cai et al. (2022) improved query propagation by employing both short- and long-term memory banks, while TransTrack (Sun et al., 2020) propagated queries across frames for continuous detection. Other extensions include P3AFormer Zhao et al. (2022), which integrates optical-flow-guided feature propagation, and CO-MOT (Yan et al., 2023), which introduces one-to-set label assignment and a COopetition Label Assignment (COLA) strategy to enhance detection performance. MeMOTR Gao & Wang (2023) further improved association by incorporating historical frame information.
MOTRv2 Zhang et al. (2023) integrated a YOLOX detector to strengthen detection with external proposals, but at the cost of additional CNN components, making it no longer a fully end-to-end solution. Moreover, its large YOLOX-L backbone (54.2M parameters) substantially increases model size compared to MOTR (40M) without fundamentally addressing MOTR’s limitations. MOTRv3 Yu et al. (2023) introduced several strategies, including Release-Fetch Supervision (RFS) to mitigate label insufficiency, Pseudo Label Distillation (PLD) from a pretrained 2D detector, and Query Grouping with Disturbance (TGD). Although effective, these designs add complexity and yield only modest improvements (e.g., +0.5 HOTA over MOTRv2). We argue that such strategies deviate from the original motivation of a simple, elegant end-to-end solution.
Importantly, none of these methods directly address the issue of query collisions between track and detection queries. We empirically demonstrate their impact in Figure 1(a), which shows the distance distribution between queries and ground-truth boxes. MATR (ours, purple) concentrates track queries near ground truth by pre-moving them before decoding, while MOTR (cyan) exhibits a more dispersed distribution, often with large deviations (mean distance 0.33 vs. 0.12 for Hungarian matching). Notably, many MOTR points cluster at $distance=1$ (IoU = 0), highlighting its difficulty in handling large motions between frames. Although MATR reduces these cases significantly, its mean distance (0.13) still slightly exceeds the ideal Hungarian matching baseline (0.12). This analysis confirms that query collisions exist and materially degrade performance: track queries risk being matched to nearby distractors, while detection queries are corrupted by noisy track predictions. Qualitative examples in Figure 5 further illustrate how such collisions affect association in crowded scenes.
Figure 2: Overview of the MATR architecture. Unlike MOTR, MATR leverages features from the Transformer Encoder (“Memory”) to update track queries $U_{\text{trk}}^{i}$ at time ii. The input frame is denoted by $I^{i}$, and the sequence of frames by $S$. $Q_{\det}$ represents learnable detection queries, and $Q_{\text{trk}}^{i}$ are the tracking results at time ii. The Motion-Aware Transformer (MAT) module is supervised by a motion loss and trained jointly with the Decoder outputs. By predicting future positions, MAT updates both track query features and positional embeddings in advance.
In summary, prior work has advanced MOT either through (1) tracking-by-detection pipelines that depend heavily on detector accuracy, or (2) query propagation strategies that aim for elegance but still suffer from fundamental limitations. While the latter remove hand-crafted association and achieve end-to-end learning, they overlook the critical issue of query collisions between track and detect queries. Our work directly addresses this gap by introducing motion-aware prediction, a simple yet effective mechanism that reduces collisions and improves consistency in end-to-end multi-object tracking.
@@ 3.1 Baseline Improvement
Since the introduction of MOTR, which was built on Deformable DETR Zhu et al. (2021), newer DETR-family frameworks such as DAB-Deformable DETR Liu et al. (2022) have demonstrated improved detection performance. To establish a stronger baseline, we adopt the bounding box propagation mechanism from DAB-DETR, where bounding boxes are propagated as positional encodings for queries. Specifically, bounding boxes are initialized randomly as $N_{\det}\times[x_{0},y_{0},h_{0},w_{0}]$, where $N_{\det}$ is the number of detection queries, and iteratively refined by successive decoder layers as $N_{\det}\times[\Delta x,\Delta y,\Delta h,\Delta w]$. It is important to note that we use only the box propagation strategy from Zhu et al. (2021), rather than adopting the full DAB model. Fully incorporating DAB increases parameters from 42M to 49M, but unexpectedly reduces HOTA from 69.4 to 67.6. We observed that larger models generated more bounding boxes, yet overall performance degraded, likely due to overfitting in the absence of additional augmentation.
Unlike MOTR, which gradually increases sequence length during training to address convergence challenges, we maintain a fixed sequence length $S$ without stability issues. For data augmentation, we simulate object entry and exit by randomly dropping track queries from previous frames, rather than artificially introducing new data. This design allows detection queries to rediscover dropped objects as newly appearing in later frames. To ensure a fair comparison with MOTR, we do not use pretrained weights from DAB-DETR. As reported in Table 1(a), this strengthened baseline surpasses MOTR by approximately 5 points, providing a solid foundation for evaluating our proposed method.
@@ 3.2 Method Overview
We propose the Motion-Aware Transformer (MAT) module, which explicitly captures the motion information of each track query by predicting its future position, thereby enhancing tracking accuracy. The overall framework is illustrated in Figure 2. Similar to many DETR-family methods, we employ an image backbone and a Deformable Transformer Encoder to extract features from the input frame $I$.
For the first frame, detection queries $Q_{\det}$ and positional embeddings PE are randomly initialized. The query tensor $Q_{\det}$ has shape $N_{\det}\times D$, where $D$ is the embedding dimension, and PE has shape $N_{\det}\times 4$, corresponding to bounding boxes $x_{0},y_{0},h_{0},w_{0}]$. A sinusoidal encoding is applied to PE to produce positional embeddings $E_{\det}$ of the same dimension as $Q_{\det}$. The sum $Q_{\det}+E_{\det}$ is passed through the Deformable Transformer Decoder to generate the initial detection results $Q_{\det}^{0}$. These detections are filtered using an IoU threshold to obtain high-quality track queries $Q_{\text{trk}}^{0}$.
@@ 3.3 Motion-aware Transformer
Unlike prior works such as MOTR Zeng et al. (2022) and MOTRv2 Zhang et al. (2023), which update track query features solely through self-attention, we argue that this design suffers from three key limitations: 1) without explicit supervision and guidance, track queries do not learn motion effectively; 2) self-attention alone fails to properly integrate the previous track queries $Q_{\text{trk}}^{i-1}$ with features from the current frame, which negatively impacts the subsequent Transformer Decoder; and 3) leaving positional embeddings unchanged can cause a mismatch between feature and positional information, further impairing training.
@@
$U_{\text{trk}}^{t-1}=Q_{\text{trk}}^{t-1}+\text{CrossAtt}(\text{SelfAtt}(Q_{\text{trk}}^{t-1}),M^{t}),$ (1)
@@
We deliberately adopt L1 loss instead of IoU-based alternatives (e.g., GIoU or DIoU) for two reasons. First, L1 loss directly penalizes deviations in both position and scale, which is crucial for synchronizing feature and positional embeddings. On the other hand, L1 has been empirically observed in DETR-family models to provide smoother optimization and more reliable convergence. By predicting and updating the full bounding box representation under this loss, we ensure that track queries remain consistent in both feature and positional spaces, preventing the Decoder from confusing trajectories across objects.
@@ 3.4 Inference Details
The inference procedure largely follows MOTR Zeng et al. (2022), with adaptations to accommodate MAT. Since ground truth is unavailable during inference, detection queries are not filtered before being processed by MAT. To handle occlusions, we filter the outputs of the Transformer Decoder. If the confidence score of a tracked object falls below the threshold $\tau_{\text{trk}}$ in the current frame, we retain its embedding temporarily as an inactive trajectory. If its confidence remains below the threshold for more than $\mathcal{T}_{\text{miss}}$ consecutive frames, the inactive trajectory is permanently removed. Similarly, detection results are discarded if their confidence falls below $\tau_{\det}$.
In all experiments, we set the thresholds to $\tau_{\det}=0.7$, $\tau_{\text{trk}}=0.5$, and $\mathcal{T}_{\text{miss}}=25$, unless otherwise specified. All other aspects of inference remain consistent with the training procedure, ensuring that MAT can be deployed seamlessly without additional heuristics.

### v2

@@ 1 Introduction
Multi-object tracking (MOT) aims to detect objects and maintain their identities across frames in a video. Recent advances in query-based detection have reshaped this task. The introduction of DETR [carion2020end] marked a turning point, as it replaced handcrafted post-processing with an end-to-end formulation. This query-based design allows the network to propagate information across frames, making end-to-end training feasible and effective.
Figure 1: Illustration of query collisions in MOTR.
@@
Building on this paradigm, TrackFormer [meinhardt2022trackformer] extended DETR to MOT by propagating detection queries forward as track queries, and concatenating them with newly initialized detection queries for subsequent frames. This formulation enabled consistent learning of track queries without manual association rules. Later, MOTR [zeng2022motr] introduced a Temporal Aggregation Network (TAN) to update track queries between frames, alleviating issues of overfitting and improving performance on challenging datasets such as DanceTrack. These advances have established query propagation as a dominant strategy in end-to-end MOT.
@@
However, a key limitation remains: existing approaches typically process detection and track queries simultaneously within a single Transformer Decoder layer. This design introduces what we call query collisions. Track queries are required to follow the same object consistently across frames, whereas detection queries are reassigned at each frame through Hungarian matching. When a track query drifts from its ground-truth location, Hungarian matching may assign it to a different object that happens to be closer, resulting in identity switches and unstable training. At the same time, detection queries suffer from noisy gradients caused by drifting track queries, further reducing association performance. Figure 1 illustrates a typical collision scenario: “A” and “B” indicate the same object at different times. In MOTR, track queries (brown boxes) must consistently track the same object (A). For newly appearing objects, Hungarian matching, similar to DETR, assigns empty detect queries (cyan boxes) to the nearest ground truth. However, track queries are not always close to their ground truth positions in the subsequent frame. As shown in the top-right figure, the red box has only 0.1 IoU with its ground truth object (A) but has a 0.7 IoU with object B, while another detect query has a 0.4 IoU with object A. Under strict Hungarian matching, the red box incorrectly assigns to object B, causing a collision. Track queries must always follow their respective objects regardless of IoU, and detect queries must not assign to already tracked object.
Figure 2: Distribution of track query distances ($1-\text{IoU}$) for different methods. We collected these distances from collision points during loss computation. Each point represents the distance between a track query’s current position and its corresponding ground truth. The dashed lines indicate the average distances. For reference, the gray dashed line shows the average distance of detect queries assigned by Hungarian matching.
@@
To overcome these limitations, we propose the Motion-Aware TransformerMATR explicitly predicts the motion of track queries across frames (Figure 1 bottom), updating both their features and positional embeddings before they enter the Transformer Decoder. By anticipating object movements, MATR reduces the gap between track queries and their ground-truth targets, thereby minimizing query collisions and aligning the training process more closely with inference behavior. This motion-aware update provides consistent optimization for both detection and tracking queries, leading to substantial gains in association accuracy without adding complex components. As shown in Figure 2, MATR (ours, purple) concentrates track queries near ground truth by pre-moving them before decoding, while MOTR (cyan) exhibits a more dispersed distribution, often with large deviations (mean distance 0.33 vs. 0.12 for Hungarian matching). Notably, many MOTR points cluster at $distance=1$ (IoU = 0), highlighting its difficulty in handling large motions between frames. Although MATR reduces these cases significantly, its mean distance (0.13) still slightly exceeds the ideal Hungarian matching (0.12). This analysis confirms that query collisions exist and materially degrade performance. MATR significantly lowers the distance between queries and their assigned objects compared to MOTR, confirming the effectiveness of motion-aware prediction. Figure 6 shows qualitative examples.
@@
In addition to introducing MATR, we strengthen the MOTR baseline with improved training strategies and updated DETR-family components such as DAB-DETR [liu2022dabdetr]. This provides a stronger foundation for comparison and ensures that the observed improvements stem from the proposed motion-aware design. Comprehensive experiments across three benchmarks validate our approach. On the DanceTrack dataset, MATR improves HOTA by more than 9 points over MOTR without additional data, and reaches a state-of-the-art HOTA of 71.3 when supplementary data are included. On SportsMOT, MATR achieves a HOTA of 72.2, establishing a new state of the art without relying on extra datasets. Finally, in multi-category tracking on BDD100k, MATR obtains 54.7 mTETA and 41.6 mHOTA, surpassing previous methods under the same training conditions.
@@
In summary, MATR addresses the issue of query collisions in DETR-style MOT by introducing explicit motion prediction. This simple yet effective design improves association accuracy across diverse scenarios, leading to state-of-the-art performance while maintaining the elegance and efficiency of an end-to-end Transformer-based tracker.
Figure 3: Overview of the MATR architecture. Unlike MOTR, MATR leverages features from the Transformer Encoder (“Memory”) to update track queries $U_{\text{trk}}^{i}$ at time ii. The input frame is denoted by $I^{i}$, and the sequence of frames by $S$. $Q_{\det}$ represents learnable detection queries, and $Q_{\text{trk}}^{i}$ are the tracking results at time ii. The Motion-Aware Transformer (MAT) module is supervised by a motion loss and trained jointly with the Decoder. By predicting future positions, MAT updates track query features and positional embeddings in advance.
@@ 2 Related Work
Tracking-by-Detection remains the most common paradigm for multi-object tracking. These methods typically adopt a two-step pipeline: first detecting bounding boxes in each frame, then associating detections across frames to construct trajectories. Performance is therefore heavily dependent on the accuracy of the detection stage. SORT [bewley2016simple] exemplifies this approach by using Hungarian algorithm [kuhn1955hungarian] for association and a Kalman filter [welch1995kalman] to predict object positions. Detections are matched to predicted boxes using intersection-over-union (IoU) scores. DeepSORT [Wojke2017simple] extends SORT by incorporating appearance features extracted by a learned embedding network, which are compared via cosine distance to improve association robustness. JDE [wang2020towards], FairMOT [zhang2021fairmot], and Unicorn [yan2022towards] jointly train detection and appearance embeddings in a unified framework. ByteTrack [zhang2022bytetrack], based on a YOLOX detector, achieves state-of-the-art performance by considering both high- and low-confidence detections. BoT-SORT [aharon2022botsort] enhances SORT by refining the Kalman filter state, introducing camera-motion compensation, and integrating Re-ID features. Recent efforts incorporate transformers into the association. TransMOT [chu2023transmot] and GTR [zhou2022global] use spatial-temporal transformers to model interactions between instances and aggregate historical context. OC-SORT [cao2023observation] improves upon SORT by re-identifying and rehabilitating lost tracks.
@@
Query Propagation has emerged as a popular alternative with the rise of DETR-family models, end-to-end approach for multi-object tracking. These methods maintain query embeddings across frames to track individual objects, eliminating the need for handcrafted post-processing. Tracktor++ [bergmann2019tracktor] pioneered this idea by leveraging an R-CNN [girshick2014rcnn] regression head to re-localize bounding boxes across frames. TrackFormer [meinhardt2022trackformer] extended Deformable DETR [zhu2021deformable] to propagate detection queries as track queries, while MOTR [zeng2022motr] introduced a Temporal Aggregation Network (TAN) to continuously update track queries. MeMOT [cai2022memot] improved query propagation by employing both short- and long-term memory banks, while TransTrack [sun2020transtrack] propagated queries across frames for continuous detection. Other extensions include P3AFormer [zhao2022tracking], which integrates optical-flow-guided feature propagation, and CO-MOT [yan2023bridging], which introduces one-to-set label assignment and a COopetition Label Assignment (COLA) strategy to enhance detection performance. MeMOTR [gao2023memotr] further improved association by incorporating historical frame information. MOTRv2 [zhang2023motrv2] integrated a YOLOX detector to strengthen detection with external proposals. MOTRv3 [yu2023motrv3] introduced RFS to mitigate label insufficiency, Pseudo Label from a pretrained 2D detector, and Query Grouping with Disturbance.
@@ 3.1 Baseline Improvement
Since the introduction of MOTR, which was built on Deformable DETR [zhu2021deformable], newer DETR-family frameworks such as DAB-Deformable DETR [liu2022dabdetr] have demonstrated improved detection performance. To establish a stronger baseline, we adopt the bounding box propagation mechanism, where bounding boxes are propagated as positional encodings for queries. Specifically, bounding boxes are initialized randomly as $N_{\det}\times[x_{0},y_{0},h_{0},w_{0}]$, where $N_{\det}$ is the number of detection queries, and iteratively refined by successive decoder layers as $N_{\det}\times[\Delta x,\Delta y,\Delta h,\Delta w]$. It is important to note that we use only the box propagation strategy from [zhu2021deformable], rather than adopting the full DAB model. Fully incorporating DAB increases parameters from 42M to 49M, but unexpectedly reduces HOTA from 69.4 to 67.6. We observed that larger models generated more bounding boxes, yet overall performance degraded, likely due to overfitting in the absence of additional augmentation.
Unlike MOTR, which gradually increases sequence length during training to address convergence challenges, we maintain a fixed sequence length $S$ without stability issues. For data augmentation, we simulate object entry and exit by randomly dropping track queries from previous frames, rather than artificially introducing new data. This design allows detection queries to rediscover dropped objects as newly appearing in later frames. To ensure a fair comparison with MOTR, we do not use pretrained weights from DAB-DETR.
@@ 3.2 Method Overview
We propose the Motion-Aware Transformer (MAT) module, which explicitly captures the motion information of each track query by predicting its future position, thereby enhancing tracking accuracy. The overall framework is illustrated in Figure 3. Similar to many DETR-family methods, we employ an image backbone and a Deformable Transformer Encoder to extract features from the input frame $I$. For the first frame, detection queries $Q_{\det}$ and positional embeddings PE are randomly initialized. The query tensor $Q_{\det}$ has shape $N_{\det}\times D$, where $D$ is the embedding dimension, and PE has shape $N_{\det}\times 4$, corresponding to bounding boxes $x_{0},y_{0},h_{0},w_{0}]$. A sinusoidal encoding is applied to PE to produce positional embeddings $E_{\det}$. The sum $Q_{\det}+E_{\det}$ is passed through the Deformable Transformer Decoder to generate the initial detection results $Q_{\det}^{0}$. These detections are filtered using an IoU threshold to obtain high-quality track queries $Q_{\text{trk}}^{0}$.
@@ 3.3 Motion-aware Transformer
Unlike prior works such as MOTR [zeng2022motr, zhang2023motrv2] which update track query features solely through self-attention, we argue that this design suffers from three key limitations: 1) without explicit supervision and guidance, track queries do not learn motion effectively; 2) self-attention alone fails to properly integrate the previous track queries $Q_{\text{trk}}^{i-1}$ with features from the current frame, which negatively impacts the subsequent Transformer Decoder; and 3) leaving positional embeddings unchanged can cause a mismatch between feature and positional information, further impairing training.
@@
$U_{\text{trk}}^{t-1}=Q_{\text{trk}}^{t-1}+\text{CrossAtt}(\text{SelfAtt}(Q_{\text{trk}}^{t-1}),M^{t})$ (1)
@@
We deliberately adopt L1 loss instead of IoU-based alternatives (e.g., GIoU or DIoU) for follow reasons. First, trajectory supervision must remain stable even when boxes have little or no overlap across frames (e.g., during fast motion or occlusion). Second, L1 loss directly penalizes deviations in both position and scale, which is crucial for synchronizing feature and positional embeddings. By predicting and updating the full bounding box representation under this loss, we ensure that track queries remain consistent in both feature and positional spaces, preventing the Decoder from confusing trajectories across objects.
@@ 3.4 Inference Details
The inference procedure largely follows MOTR [zeng2022motr], with adaptations to accommodate MAT. Since ground truth is unavailable during inference, detection queries are not filtered before being processed by MAT. To handle occlusions, we filter the outputs of the Transformer Decoder. If the confidence score of a tracked object falls below the threshold $\tau_{\text{trk}}$ in the current frame, we retain its embedding temporarily as an inactive trajectory. If its confidence remains below the threshold for more than $\mathcal{T}_{\text{miss}}$ consecutive frames, the inactive trajectory is permanently removed. Similarly, detection results are discarded if their confidence falls below $\tau_{\det}$. In all experiments, we set the thresholds to $\tau_{\det}=0.7$, $\tau_{\text{trk}}=0.5$, and $\mathcal{T}_{\text{miss}}=25$, unless otherwise specified.